OpenStudy (anonymous):

write this in trinomial form 2b^2-11-21

OpenStudy (mertsj):

Well -11-21 = -32 so it is NOT a trinonial

OpenStudy (anonymous):

really its in my test

OpenStudy (mertsj):

Well then, you have posted something incorrectly

OpenStudy (anonymous):

how abt in factored form ..

OpenStudy (anonymous):

is it 2b^2-11b-21?

OpenStudy (mertsj):

Why don't you just post it correctly?

OpenStudy (anonymous):

did you type it incorrectly?

OpenStudy (anonymous):

noo :( it was,, write this trinomial in factored form

OpenStudy (mertsj):

You left out the b

OpenStudy (mertsj):

You have the b^2 but you left out the b. If you want help, post it correctly.

OpenStudy (anonymous):

yeah :( sorry am nervious

OpenStudy (mertsj):

Well now we know that you are nervous and sorry. Please post the question correctly.

OpenStudy (anonymous):

but can you help me ???

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the problem is factor 2b^2-11b-21, am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to factor things like this?

OpenStudy (anonymous):

no I have problem understanding them

OpenStudy (anonymous):

um... I don't quite know how to explain the way I do it... so you'll have to get someone else to help sorry I hope you get your answer and merry christmas

OpenStudy (anonymous):

well thanks any ways .. and you too happy merry Christmas

Directrix (directrix):

@monchiiz68 Factor by grouping: 2b² -11 b -21 First, look for two numbers that multiply to (-21)*(2) AND also add to -11. That would be 3 and -14 2b² -11 b -21 = 2b² - 14b + 3b - 21 = 2b ( b - 7 ) + 3( b - 7) = [Note that (b -7) is a common factor, so] (b -7) (2b + 3)